OpenStudy (anonymous):

Find the solution of the exponential equation, rounded to four decimal places. 5^x=4^(x+1)

OpenStudy (amistre64):

5^x=4^(x+1) 5^x=4^x * 4 5^x/4^x = 4 (5/4)^x = 4 mightbe hard to do without a calculator or some higher methods

OpenStudy (amistre64):

(5/4)^x = 4 x ln(5/4) = ln(4) x = ln(4)/ln(5/4) x = ln(4)/ [ln(5) - ln(4) ]

OpenStudy (anonymous):

thank you!!!